A 0.580 g sample of a carboxylic acid is burned in oxygen, producing 1.03 g of CO2 and 0.426 g of H2O. Determine the empirical formula of the carboxylic acid.
Answer #1

Answers Given data Total organic compound -0.5809 co mass = 1.039 uu H₂O mass = 0.426 g 6.46 = 12 x Massotro, X100 Total Orga90 = 100-%(-494 = 100-480432-8.160%. = 100 - 56-592 0%. = 43.408.7 Empirical formula o улу Сло element si element % mass mass